OpenStudy (waheguru):

How do you find the equation of a line when you are only given slope and the x -intercept

OpenStudy (anonymous):

y=mx+b!

OpenStudy (waheguru):

yea but how do i find b

OpenStudy (waheguru):

or y

OpenStudy (anonymous):

m being slope, and you have a point so solve for b, the x-intercept being your point

OpenStudy (anonymous):

x intercept means you have the point (a,0) use point - slope \[y-0=m(x-a)\]

OpenStudy (anonymous):

confused

OpenStudy (mertsj):

if you know the x intercept you know a point. Use point slope form.

OpenStudy (anonymous):

for example if the intercept was (1,0) then 0= m(1) +b

OpenStudy (anonymous):

okay i listenin

OpenStudy (waheguru):

ok the slope is -1/2 and x int is -10 so what do i do

OpenStudy (anonymous):

you would divide or subtract i think

OpenStudy (waheguru):

help

OpenStudy (anonymous):

huh